# Measuring the Dunkelflaute: How (not) to analyze variable renewable energy shortage

### v1

@@ 3.3 Comparability, scope of analysis, and critical appraisal
Implicit comparability. Unlike VRE droughts, analyzing PRL events does not require pre-processing, such as a FLH adjustment, to enable comparability. PRL events are denoted in energy (e.g., megawatt hours), which is readily comparable across different technological, spatial, and temporal scopes.
@@
Setting: Exogenous residual load. The analysis of residual load time series can be approached in two ways. On option is to use exogenous assumptions for both VRE availability factors and demand time series. This setup facilitates the exploration of various user-specific VRE capacity mix scenarios.
@@
Setting: Optimized residual load. Alternatively, residual load time series can be retrieved from a scenario optimized by an energy system model. This method is more resource-intensive due to the need for comprehensive input data and complex modeling assumptions, but also less dependent on exogenous capacity assumptions. It further enables an analysis of policy-relevant scenarios, such as those outlined in strategic documents like the Ten Year Network Development Plan for the European electricity infrastructure entso-e_ten_2022 .
[…]
@@
General Limitation. During an extreme PRL events, brief periods with very high VRE generation requiring partial curtailment may occur. None of the discussed identification methods is able to account for such curtailed energy, which can lead to underestimating the energy deficit and duration of extreme PRL events ruhnau_storage_2022 . More importantly, insights from PRL event analysis are only applicable to specific assumptions on the VRE portfolio and load, or driven by the parameterization of optimizing models. Therefore, when discussing characteristics such as the duration, frequency, or energy deficit of PRL events, contextualization is essential. This is particularly important for a meaningful comparison of the findings of different studies, and for communication with policy-makers or the general public.
@@
Limitations: exogenous residual load. By design, an exogenously derived residual load time series does not account for electricity exchange and potentially neglects flexibility from conventional demand and electrified loads from the heating, transport, or industry sector. Accurately estimating the operation of these loads without the use of an energy system model is challenging and may require numerous assumptions.
@@
Limitations: optimized residual load. As for residual load time series generated by optimizing energy system models, model outcomes rely on the used modeling framework and its underlying optimization rationale, the level and detail of technology representation, and techno-economical assumptions, such as available technologies, costs, meteorological resources gils_modeling_2022 ; gils_model-related_2022 ; van_ouwerkerk_impacts_2022 . For instance, ceteris paribus, changes in storage investment costs can significantly affect storage investment and usage in a modeled scenario, thus altering the energy deficit or duration of obtained PRL events.
Note that an optimized residual load resulting from energy system modeling already takes into account the optimal capacities and dispatch decisions of electricity storage and other flexibility options. This inherently affects PRL event characteristics.
@@ 4 Conclusion and recommendations
4.1 Contribution
Gap in literature & suggested terminology. VRE shortage has drawn research interest from various fields, including hydrology, wind and solar energy, energy system modeling, or climatology, and has been analyzed using a range of methods and disparate naming. Yet, there is no structured overview and explanation of these methods using harmonized terminology. It thus remains unclear as to what methodological approach is useful for different settings of VRE shortage analysis, which can be studied based on time series of VRE availability or positive residual load. For clarity, we suggest “variable renewable energy shortage” as a generic term covering both “variable renewable droughts” as periods with low resource availability of a single or multiple VRE technologies and “positive residual load events” as periods with a net imbalance of electric load and VRE supply.
Contribution. This paper addresses this gap by providing a methodological overview and guidance for defining and quantifying VRE shortage events across technologies and regions, discussing different approaches for both VRE availability and PRL time series. While not claiming to be exhaustive, we categorize and discuss relevant existing shortage event definitions and identification methods, propose methodological improvements, and elaborate on how to enable compatibility in cross-technology or multi-regional analyses. Our work seeks to support the harmonization of terminology across related fields. We also aim to aid in advancing future research on VRE shortage events by facilitating a better interpretation of such studies’ potential findings.
@@ 4.2 What (not) to do in variable renewable energy shortage analysis
Good practices. We identify a range of good practices for VRE shortage analysis to obtain accurate, broadly applicable, and easy-to-interpret findings on the duration, frequency, and energy deficit of VRE shortage events. This is important for guiding policy or infrastructure decisions for future energy systems with high VRE penetration.
@@
Good practices: specify method. We advocate for future work to clearly and unambiguously specify the shortage period definitions and identification methods used, as well as key parameter choices such as thresholds or averaging intervals. Methods that are poorly defined or not transparent hinder the accurate interpretation and contextualization of findings, in both research and policy discourse.
@@
Good practice: methodical requirements. For accurate identification, VRE shortage analysis should meet a range of methodological criteria. First, VRE shortage periods should be defined according to the event perspective, i.e., capturing the full extent of qualifying periods allowing for event duration ranging from just a few time steps to extensive weeks or even months. Second, identified shortage events should be unique, avoiding double counting and overlap with adjacent events. Third, the method should allow for pooling of closely occurring events to include periods that independently may not qualify as VRE drought or PRL events but are adjacent to periods of low availability or high residual load.
@@
Good practice: recommended VRE shortage characteristics. In general, reporting the VRE shortage duration appears to be equally relevant to VRE drought and PRL analyses. Yet, claims on identified durations have to be contextualized, as they depend on the identification method used. Further, evaluating energy deficits of PRL events is valuable for understanding the demand for firm generation or long-duration storage energy in renewable power sectors. Conversely, even tho this metric is commonly applied in VRE drought analysis across different bodies of literature, it appears to be less relevant in such settings, as it strongly depends on the threshold parameter choice and cannot be associated with a power sector attribute.
[…]
@@
Good practice: comparative analyses. Analyzing VRE droughts across multiple regions and/or for different VRE technologies or technology portfolios poses comparability challenges. It is essential to base the analysis on uniform average generation (or generation potential) across compared systems to avoid a FLH bias in obtained drought characteristics. Threshold scaling for each investigated system according to a fraction of the mean capacity factors is an effective option for enabling comparability.
@@
Good practice: threshold selection. The selection of the respective fraction remains a sensitive parameter choice and requires careful consideration. Too high thresholds may identify long-lasting periods with below-average VRE availability rather than actual VRE droughts. Conversely, very low thresholds may find very severe but potentially brief and isolated droughts. We advise exploring various threshold values between these extremes, ranging from near-zero to the average availability factor. This ensures a comprehensive capture of the full spectrum of drought events, from mild to very severe.
@@
Good practices: open code and data. For transparency and reproducibility, it further appears vital to document and provide the underlying data and the algorithms used, adhering to open-source and open-data principles pfenninger_opening_2018 . Such practices will not only foster transparency and replicability, but also strengthen the overall accuracy and practical usefulness of VRE drought analyses.
@@
Bad practices. Based on our analysis, we identify a range of bad practices for VRE shortage analysis. Searching for fixed-duration shortage windows underestimates shortage duration while inflating the number of identified shortage periods. We thus strongly recommend not using the window perspective. Methods yielding ambiguous sets of events that may change for varying method parameterization, such as FMBT and FMAZ, should be avoided due to the limited applicability of obtained shortage characteristics. Next, we advise against the use of methods incapable of pooling mutually dependent shortage events, such as CBT and CAZ, as obtained results overestimate the number of events while underestimating event duration and energy deficit. Finally, in VRE drought studies, using absolute thresholds can further lead to misleading comparisons based on normalized generation capacities rather than actual energy production, resulting in potentially flawed conclusions.
@@ 4.3 Methodological trade-offs and outlook
Trade-offs between VRE drought and PRL analysis. Renewable energy shortage can be studied based on time series of VRE availability or positive residual load. While both approaches have their merits, a trade-off between policy relevance and the universality of insights emerges. VRE drought analysis based on VRE availability data typically requires fewer assumptions, yielding conclusions with broader applicability. PRL events analyses aim to be more policy-oriented, as they quantify the mismatch between electricity demand and VRE supply, which is especially relevant for highly renewable decarbonization scenarios. However, the latter approach requires extensive assumptions on the composition of the renewable generation portfolio and demand-side flexibility. This may necessitate detailed energy system modeling, whose outcomes hinge upon underlying optimization rationales, technology representation, and parameterization. This potentially limits the validity of PRL analyses to the particular scenarios investigated.
Trade-offs between single and multiple technologies VRE drought analysis. Such a trade-off also applies to the technological scope of VRE drought analysis. Aside from methodological assumptions inherent to the generation of VRE availability time series, e.g., such for reanalysis, insights from single-technology analysis are broadly applicable but less policy-relevant for future energy systems that rely on a mix of VRE technologies. Conversely, analyzing a VRE portfolio is more policy-oriented but involves energy modeling or exogenous assumptions on the composition of the investigated capacity mix. VRE capacity mixes determined by an energy modeling are optimal only in the context of the other optimized system components and flexibility options that are represented in the model (e.g., storage, electricity exchange, or direct-electrified applications from the heat, industry, or transport sector). Similarly, exogenous capacity mixes rely on user-specific assumptions. These aspects limit the validity of insights to investigated scenarios.
Outlook. While we provide general guidance on the advantages and drawbacks of different VRE shortage identification methods, a thorough quantitative comparison would be desirable, ideally for a range of VRE technologies and electric load profiles for different world regions and weather years. Another key aspect that requires attention is the choice of the threshold parameter, particularly in distinguishing between severe VRE droughts and prolonged periods of below-average availability. This distinction could support the selection of relevant weather years to enable weather- and climate-resilient energy system analyses and well-informed decarbonization policies. Additionally, assessing the sensitivity of outcomes to different VRE availability datasets would be useful and could be valuable feedback to the meteorological research community that provides and continuously improves such datasets craig_overcoming_2022 .

### v2

@@ 3.3 Comparability, scope of analysis, and critical appraisal
Unlike VRE droughts, analyzing PRL events does not require pre-processing, such as a FLH adjustment, to enable comparability. PRL events are denoted in energy (e.g., megawatt hours), which is readily comparable across different technological, spatial, and temporal scopes.
@@
The analysis of residual load time series can be approached in two ways. One option is to use exogenous assumptions for both VRE availability factors and demand time series. This setup facilitates the exploration of various user-specific VRE capacity mix scenarios.
@@
Alternatively, residual load time series can be retrieved from a scenario optimized by an energy system model. This method is more resource-intensive due to the need for comprehensive input data and complex modeling assumptions, but also less dependent on exogenous capacity assumptions. It further enables an analysis of policy-relevant scenarios, such as those outlined in strategic documents like the Ten Year Network Development Plan for the European electricity infrastructure [68].
Characterizing extreme PRL events, i.e., their duration, energy deficit, frequency, etc., allows exploring the need for power sector flexibility, which can be provided by long-duration storage or other flexibility options. The SPAadj, accounting for energy losses of such flexibility options, appears to be particularly useful in this respect.
@@
During an extreme PRL events, brief periods with very high VRE generation requiring partial curtailment may occur. None of the discussed identification methods is able to account for such curtailed energy, which can lead to underestimating the energy deficit and duration of extreme PRL events [22]. More importantly, insights from PRL event analysis are only applicable to specific assumptions on the VRE portfolio and load, or driven by the parameterization of optimizing models. Therefore, when discussing characteristics such as the duration, frequency, or energy deficit of PRL events, contextualization is essential. This is particularly important for a meaningful comparison of the findings of different studies and for communication with policy-makers or the general public.
@@
By design, an exogenously derived residual load time series does not account for electricity exchange and potentially neglects flexibility from conventional demand and electrified loads from the heating, transport, or industry sector. Accurately estimating the operation of these loads without the use of an energy system model is challenging and may require numerous assumptions.
@@
As for residual load time series generated by optimizing energy system models, model outcomes rely on the used modeling framework and its underlying optimization rationale, the level and detail of technology representation, and techno-economical assumptions, such as available technologies, costs, meteorological resources [69, 70, 71]. For instance, ceteris paribus, changes in storage investment costs can significantly affect storage investment and usage in a modeled scenario, thus altering the energy deficit or duration of obtained PRL events. Further, an optimized residual load already takes into account the endogenously determined investment and dispatch decisions of electricity storage and other flexibility options. This inherently affects PRL event characteristics.
@@ 4 Conclusion and recommendations
4.1 Summary
VRE shortage has drawn research interest from various fields, including hydrology, wind and solar energy, energy system modeling, or climatology, and has been analyzed using a range of methods and disparate naming. Yet, there is no structured overview and explanation of these methods using harmonized terminology. It thus remains unclear as to what methodological approach is useful for different settings of VRE shortage analysis, which can be studied based on time series of VRE availability or positive residual load. For clarity, we suggest “variable renewable energy shortage” as a generic term covering both “variable renewable droughts” as periods with low resource availability of a single or multiple VRE technologies and “positive residual load events” as periods with a net imbalance of electric load and VRE supply.
This paper addresses this gap by providing a methodological overview and guidance for defining and quantifying VRE shortage events across technologies and regions, discussing different approaches for both VRE availability and PRL time series. While not claiming to be exhaustive, we categorize and discuss relevant existing shortage event definitions and identification methods, propose methodological improvements, and elaborate on how to enable compatibility in cross-technology or multi-regional analyses. Our work seeks to support the harmonization of terminology across related fields. We also aim to aid in advancing future research on VRE shortage events by facilitating a better interpretation of such studies’ potential findings.
@@ 4.2 What (not) to do in variable renewable energy shortage analysis
We identify a range of good practices for VRE shortage analysis to obtain accurate, broadly applicable, and easy-to-interpret findings on the duration, frequency, and energy deficit of VRE shortage events. This is important for guiding policy or infrastructure decisions for future energy systems with high VRE penetration.
@@
We advocate for future work to clearly and unambiguously specify the shortage period definitions and identification methods used, as well as key parameter choices such as thresholds or averaging intervals. Methods that are poorly defined or not transparent hinder the accurate interpretation and contextualization of findings, in both research and policy discourse.
@@
For accurate identification, VRE shortage analysis should meet a range of methodological criteria. First, VRE shortage periods should be defined according to the event perspective, i.e., capturing the full extent of qualifying periods allowing for event duration ranging from just a few time steps to extensive weeks or even months. Second, identified shortage events should be unique, avoiding double counting and overlap with adjacent events. Third, the method should allow for pooling of closely occurring events to include periods that independently may not qualify as VRE drought or PRL events but are adjacent to periods of low availability or high residual load.
@@
In general, reporting the VRE shortage duration appears to be equally relevant to VRE drought and PRL analyses. Yet, claims on identified durations have to be contextualized, as they depend on the identification method used. Further, evaluating energy deficits of PRL events is valuable for understanding the demand for firm generation or long-duration storage energy in renewable power sectors. Conversely, even though this metric is commonly applied in VRE drought analysis across different bodies of literature, it appears to be less relevant in such settings, as it strongly depends on the threshold parameter choice and cannot be associated with a power sector attribute.
For VRE drought analysis, the VMBT and SPA methods align well with these methodological requirements. The VMBT method finds events with the longest duration but lacks meaningful energy deficits, which is a reasonable cost for accurate duration identification. In the case of PRL analyses, we recommend the use of SPA or its adjusted version, SPAadj. Besides adhering to the outlined criteria, these methods allow for accurately quantifying both duration and the energy deficit of PRL events.
@@
Analyzing VRE droughts across multiple regions and/or for different renewable technologies or technology portfolios poses comparability challenges. It is essential to base the analysis on uniform average generation (or generation potential) across compared systems to avoid a full-load hour bias in obtained drought characteristics. Threshold scaling for each investigated system according to a fraction of the mean availability factors is an effective option for enabling comparability.
@@
The selection of the respective fraction remains a sensitive parameter choice and requires careful consideration. Too high thresholds may identify long-lasting periods with below-average renewableavailability rather than actual VRE droughts. Conversely, very low thresholds may find very severe but potentially brief and isolated droughts. We advise exploring various threshold values between these extremes, ranging from near-zero to the average availability factor. This ensures a comprehensive capture of the full spectrum of drought events, from mild to very severe.
@@
For transparency and reproducibility, it further appears vital to document and provide the underlying data and the algorithms used, adhering to open-source and open-data principles [72]. Such practices will not only foster transparency and replicability but also strengthen the overall accuracy and practical usefulness of VRE drought analyses.
@@
Based on our analysis, we identify a range of bad practices for variable renewable energy shortage analysis. Searching for fixed-duration shortage windows underestimates shortage duration while inflating the number of identified shortage periods. We thus strongly recommend not using the window perspective. Methods yielding ambiguous sets of events that may change for varying method parameterization, such as FMBT and FMAZ, should be avoided due to the limited applicability of obtained shortage characteristics. Next, we advise against the use of methods incapable of pooling mutually dependent shortage events, such as CBT and CAZ, as obtained results overestimate the number of events while underestimating event duration and energy deficit. Finally, in VRE drought studies, using absolute thresholds can further lead to misleading comparisons based on normalized generation capacities rather than actual energy production, resulting in potentially flawed conclusions.
@@ 4.3 Methodological trade-offs and outlook
Renewable energy shortage can be studied based on time series of VRE availability or positive residual load. While both approaches have their merits, a trade-off between policy relevance and the universality of insights emerges. Drought analyses based on VRE availability data typically require fewer assumptions, yielding conclusions with broader applicability. PRL events analyses aim to be more policy-oriented, as they quantify the mismatch between electricity demand and variable renewable supply, which is especially relevant for highly renewable decarbonization scenarios. However, the latter approach requires extensive assumptions on the composition of the renewable generation portfolio and demand-side flexibility. This may necessitate detailed energy system modeling, whose outcomes hinge upon underlying optimization rationales, technology representation, and parameterization. This potentially limits the validity of PRL analyses to the particular scenarios investigated.
Such a trade-off also applies to the technological scope of VRE drought analysis. Aside from methodological assumptions inherent to the generation of variable renewable availability time series, e.g., such for reanalysis, insights from single-technology analysis are broadly applicable but less policy-relevant for future energy systems that rely on a mix of renewable technologies. Conversely, analyzing a VRE portfolio is more policy-oriented but involves energy modeling or exogenous assumptions on the composition of the investigated capacity mix. VRE capacity mixes determined by an energy modeling are optimal only in the context of the other optimized system components and flexibility options that are represented in the model (e.g., storage, electricity exchange, or direct-electrified applications from the heat, industry, or transport sector). Similarly, exogenous capacity mixes rely on user-specific assumptions. These aspects limit the validity of insights to investigated scenarios.
While we provide general guidance on the advantages and drawbacks of different VRE shortage identification methods, a thorough quantitative comparison would be desirable, ideally for a range of variable renewable technologies and electric load profiles for different world regions and weather years. Another key aspect that requires attention is the choice of the threshold parameter, particularly in distinguishing between severe VRE droughts and prolonged periods of below-average availability. This distinction could support the selection of relevant weather years to enable weather- and climate-resilient energy system analyses and well-informed decarbonization policies. Additionally, assessing the sensitivity of outcomes to different VRE availability datasets would be useful and could be valuable feedback to the meteorological research community that provides and continuously improves such datasets [50].